# TALES: A Taxonomy and Analysis of Cultural Representations in LLM-generated Stories

### v1

@@ 5.1. Question Bank Creation (TALES-QA)
We used the misrepresentation spans identified by our expert annotators to generate questions. Specifically, GPT $4.1$ was instructed to formulate a single standalone question given the story text, the annotated span, its categorization and rationale as inputs. The model was instructed through prompting to ensure that the resulting candidate question directly addressed the contextual inconsistency, was interpretable without access to the original story, and avoided ambiguity. The questions were produced in multiple formats, including one-word-answer, fill-in-the-blank, multiple-choice, and true-or-false questions. The generated questions were in the same language as the story they were derived from, which gave us questions in both English and other $13$ Indic languages.
@@
Next, all generated questions were reviewed by human experts and refined whenever needed. We worked with $61$ participants. Of these, $44$ were language and culture experts from AI4Bharat who verified questions for the languages of their expertise. The remaining $17$ participants were recruited from our institution and reviewed English questions for the regions they had lived in. Before beginning the task, all participants were provided with clear written instructions outlining the evaluation criteria, available in Appendix G.
Each question was independently assessed along four dimensions using a yes/no format: (1) validity: whether the question was valid; (2) uniqueness: whether it admitted a single unambiguous answer; (3) cultural grounding: verifying the presence of a cultural element; and (4) answer accuracy: confirming that the expected response was correct. In cases where a question did not satisfy one or more of these criteria, participants were instructed to refine it to meet the required standards. If the question received a ‘no’ on any of the dimensions and the verifier did not suggest a refinement, the question was excluded from TALES-QA. This protocol ensured that the final set of questions were clear, accurate, and culturally appropriate.
The question generation and verification process resulted in a TALES-QA, a question bank comprising $5$ different types of questions: one-word answer, one-phrase answer, fill-in-the-blank, multiple-choice, and true/false. The final bank contained a total of 1,683 questions, with 1,115 in Indic languages and 568 in English (see Table 8). The majority of the questions are of one-word type, which are particularly effective at testing cultural knowledge. Unlike True/False and MCQ formats, one-word questions offer a more robust means of assessment because they minimize the possibility of random guessing.
5.2. Evaluation Methodology
We evaluated all $6$ models on TALES-QA to assess their cultural knowledge. To prevent models from producing long-form responses, we provided instructions tailored to each question type. To mitigate potential false negatives arising from Unicode or script matching issues in Indic languages, and following the recent success of auto raters for answer matching (Chandak et al., 2025), we used GPT 44o as a judge to compare each model’s responses against the correct answers (prompt provided in Appendix H). Additionally, to capture variability and uncertainty in model outputs, we sampled five responses from the model being tested and computed average across all of them. We measured each model’s accuracy on the entire question bank. To further investigate whether misrepresentations were primarily due to a lack of relevant cultural knowledge, we also measured each model’s accuracy on the subset of questions derived specifically from misrepresentations present in the stories generated from that same model.
Aya32B
Qwen3-32B
Gemma3-27B
GPT4.1
LLaMA3.3
Gemini2.5Pro
@@ 5.2. Evaluation Methodology
Table 9. Accuracy of six models in answering cultural knowledge questions on the full TALES-QA question bank (full dataset) and on questions derived from misrepresentations generated by that model (model subset). Best performance in each row is bold
@@ 5.3. Findings
We also observed that for the closed-source models, GPT $4.1$ and Gemini 2.52.5Pro, accuracy decreased on the model-specific subsets compared to their performance on the overall English question bank. In contrast, for the open-source models, Gemma 33 2727B and LLaMa $3.3$ 7070B, accuracy increased on the model-specific subsets. This suggests that the questions derived from closed-source models may be more challenging than those derived from open-source models. Moreover, even for the best models, there exists headroom for improvement (from 86.386.3%) in answering cultural knowledge questions, in both English and Indic languages. We hope that TALES-QA will be a useful resource for testing cultural knowledge from India, in English and $13$ Indic languages, for future models.
@@ 6.1. Community-Centred Evaluations of Cultural Representation
Evaluation of an AI system, for any capability, requires a number of assumptions and design decisions (Jacobs and Wallach, 2021; Wallach et al., 2024). Such decisions include, albeit non-exhaustively, (1) what should be evaluated? (2) what is the desirable outcome? and (3) how it should be evaluated? These assumptions and decisions are often implicit, almost exclusively made by developers of AI systems, and reflect the priorities and incentive structures embedded within AI development. Yet work has shown that values prioritized by model developers and the general public may not align (Jakesch et al., 2022) and what different communities consider desirable is subjective (Diaz and Madaio, 2023). This issue is likely to be exacerbated when AI developers do not represent the heterogeneity of the users and disproportionately reflect privileged sociocultural identities (Sambasivan et al., 2021). For evaluating cultural representation, such misalignment can be problematic because needs and desires of cultural representations are deeply intertwined with lived experiences of communities. It is thus essential to account for the needs and priorities of these communities because the evaluation decisions directly impact, and can potentially harm these communities.
Keeping this in mind, we ground our evaluation of cultural misrepresentations by LLMs in the deep cultural understanding of community members. From the outset, we engage with participants with diverse cultural identities and lived experiences from across India to create a taxonomy of cultural misrepresentations in LLM-generated stories. We then use this taxonomy for a larger-scale human evaluation of 6 LLMs in 14 languages by 108 expert annotators who come from 71 different regions in India. This allowed us to conduct fine-grained analyses of cultural misrepresentations leading to valuable insights that illuminate future research directions, including: (1) gaps in multilingual capabilities and cultural representation for peri-urban regions, (2) highlighting the cultural entities (practices, norms, and food) that are often misrepresented, and (3) demonstrating the gap between models’ cultural knowledge in answering standalone questions and ability to use this knowledge in generative settings. Finally, the taxonomy of cultural misrepresentations (TALES-Tax) and questions of cultural knowledge for diverse Indian identities (TALES-QA), will be useful for future research in this area.
@@ 6.2. Minoritized Communities within Minoritized Communities
Despite AI systems being deployed globally, they continue to better represent and serve some identities (Durmus et al., 2023; AlKhamissi et al., 2024). But, the studies of harms or cultural competence are often presented as a dichotomy between “western” and “non-western” contexts (Qadri et al., 2025). However, this formulation presumes a monolithic cultural identity of the minoritized “non-western” region. Treating the diverse cultural identities within any (minoritized) region, including India, as a homogenous group is likely to prevent a finer-grained understanding of how AI systems may (mis)represent or harm minoritized communities within this minoritized region. India is home to people from diverse linguistic, ethnic, religious, socioeconomic, and educational backgrounds. The axes of disparities among these communities are distinct from the rest of the world (Bhatt et al., 2022). Prior work has found that biases along these unique axes, such as caste, disability, religion, and region within India are reinforced by AI systems (Bhatt et al., 2022; Khandelwal et al., 2024; Phutane and Vashistha, 2025) and are difficult to overcome (Seth et al., 2025).
Our work echoes and adds to these findings: cultural misrepresentations by LLMs in the Indian context are not only prevalent, but they are exacerbated in lower-resourced Indic languages and for identities from less well-known regions (§4.3). This has particularly concerning implications for Indian users of digital and AI technology. As of 2021, India has the second-highest number of active internet users in the world, accounting for over 55% of its population with widespread penetration across both rural and urban areas (World Bank, 2025a; Wikipedia contributors, 2025). At the same time, over 60% of the population resides in rural areas (World Bank, 2025b) and 50% of Indian users prefer to access internet in Indic languages (AFPR, 2025; Statista, 2025). Presumably, a significant proportion of these users use AI, or at least encounter AI generations, facing unforeseen harms. This underscores the urgent need for building evaluations of cultural representations that engage with the lived experiences of diverse communities.
We take a step in this direction by zooming in on the Indian context and engaging with participants from a wide variety of regional, linguistic, and sociocultural backgrounds for evaluating cultural misrepresentations by LLMs. We acknowledge that building language technologies in the Indian context has received attention on various fronts, like curating pre-training corpora (Khan et al., 2024), developing benchmarks (Verma et al., 2025; Singh et al., 2024; Dev et al., 2023; Seth et al., 2024), building foundation models (Kallappa et al., 2025; Choudhury et al., 2025), evaluating fairness (Bhatt et al., 2022; Khandelwal et al., 2024) and cultural competence (Maji et al., 2025). We hope to see this trend continue and benefit the minoritized sociocultural communities within India.
@@ 6.3. Designing Systems for Failure Modes
As AI systems get deployed in public-facing technologies, it is understandable that they may not have the ability to satisfactorily fulfil a user’s request. In the context of cultural representation, a system may encounter a user’s request requiring it to represent a culture that is beyond its current capabilities. Among other reasons, this could happen because the user’s request is ambiguous, the model does not have the requisite parametric knowledge, or there is more than one ways to fulfil the request. Unless the system is specifically designed to handle such situations, it may fall in one of two failure modes: (1) It generates generic text which does not include any specific cultural elements (no-representation), or (2) hallucinates information about the culture and generates text that is rife with inaccuracies around cultural elements (misrepresentations). Based on our findings (§4.3), while Llama3.3-70B generated fewer misrepresentations per story, it also had low CSI counts in the stories to begin with. This indicates that Llama3.3-70B might be generating generic stories that do not incorporate enough cultural nuance. On the other hand, Qwen3-32B and Aya-32B generate high number of CSIs in the stories but also have a larger number of misrepresentations per story. In other words, we could think of Llama3.3-70B being prone to the failure mode of no-representation, while Qwen3-32B and Aya-32B as being prone to the failure mode of misrepresentation.
Neither misrepresentation nor no-representation is an ideal outcome. In fact, the participants we engaged with indicated that the result of both these failure modes is—be it inaccuracies in cultural elements or oversimplification of cultural elements—dissatisfying. Thus, we argue that AI systems must be thoughtfully designed to handle situations where the system encounters requests that may be underspecified, subjective, cannot be fulfilled based on the parametric or external knowledge that the system can access.
AI systems may be designed to handle such situations in a variety of ways to prevent giving outrightly wrong or misleading answers. For example, the LLM may be designed to ask follow-up questions (Li et al., 2024a, 2025), abstain from answering (Wen et al., 2025; Brahman et al., 2024), or provide multiple possible outputs (Sorensen et al., 2024). Each of these design decisions may impact the user’s experience in different ways and the best way to handle such requests is contextual, and remains an open area of research. Moreover, such a design must also account for the sociocultural context of the deployment, since users across cultures may have varying expectations (Fortunati et al., 2022; Ge et al., 2024) and in turn varying tolerance to different model behaviours (Lucy et al., 2023). As AI systems become integrated into people’s daily lives, future research must investigate the impacts of their design or behaviour on user experiences and the harms they may face.
@@ 7. Limitations and Future Work
Although we view this work as an important step towards understanding and evaluating misrepresentation in generated stories, we acknowledge a few important limitations of our work. First, all focus groups and surveys were conducted on English-language stories generated by GPT-$4$, primarily to facilitate consistent interpretation and discussion within the research team. Further research is required to assess whether any new categories would emerge from other languages or from other models. Second, our evaluation study focused primarily on cultural contexts within India, and thus the findings were shaped by the perspectives most familiar to our expert annotators. While this still provides valuable insights, future work could expand data collection beyond India to capture a broader range of cultural contexts and to examine how similar misrepresentations manifest globally. Third, while our analysis captured localized misrepresentations by focusing on cultural nuance at the city-, town-, and village-level, it also introduced challenges to measure agreement, as finding sufficient expert annotators familiar with the same region was difficult.
Our work also points to concrete future research directions. To begin with, while our taxonomy provides a structured way to identify and analyze misrepresentations, future work should build on this to develop concrete strategies for mitigating the cultural misrepresentations we observed. Finally, one particularly interesting direction of future work is to investigate factors that explain why LLMs misrepresent cultures in open-ended generations despite possessing the requisite cultural knowledge.
@@ 8. Conclusion
Acknowledgments
We are grateful to the participants of our focus groups and survey, and the expert annotators of our study for their time, expertise, and thoughtful engagement with our study. We thank folks at AI4Bharat, specifically, Mitesh Khapra and Sounak Dutta, for facilitating the annotation process. This work was supported in part by the AI2050 program at Schmidt Sciences (Grant G-24-66186) and a grant from Google Research (thanks to Dinesh Tewari). We are grateful to Kinshuk Vasisht for his feedback and for being the external coder for the development of our taxonomy. We thank Jimin Mun, Akhila Yerukola, Joel Mire, Mansi Gupta, Lora Aroyo, Partha Talukdar and Deepak Varuvel Dennison for their feedback on the manuscript. Lastly, we are grateful to Vani A for her help with project coordination and administration.
@@ Appendix A Demographic details for focus groups and surveys
Please refer Table 10
Focus Group 1: Jodhpur, Ahmedabad, Jaipur (Gujarat and Rajasthan)
Focus Group 2: Kannur, Mayiladuthurai, Chennai (Kerala and Tamil Nadu)
Focus Group 3: Bhubaneswar, Kolkata, Bargarh (Odisha and West Bengal)
@@ Appendix B Questions Used by Facilitator in Focus Groups and Surveys
      What do they look for in a story, is relatability important?
@@
  • Questions after participants have completed reading the stories:
@@
      If they read multiple stories: are there any similarities between them?

### v2

@@ 5.1. Question Bank Creation (TALES-QA)
We used the misrepresentation spans identified by our expert annotators to generate questions. Specifically, GPT $4.1$ was instructed to formulate a single standalone question given the story text, the annotated span, its categorization and expert annotator’s rationale as inputs. The model was instructed through prompting to ensure that the resulting candidate question directly addressed the contextual inconsistency, was interpretable without access to the original story, and avoided ambiguity. The questions were produced in multiple formats, including one-word-answer, fill-in-the-blank, multiple-choice, and true-or-false questions. The generated questions were in the same language as the story they were derived from, which gave us questions in both English and other $13$ Indic languages.
@@
Next, all generated questions were reviewed by human experts and refined whenever needed. We worked with $61$ participants. Of these, $44$ were language and culture experts from AI4Bharat, the AI organization that we partnered with for the large-scale annotation effort (§4.1). These language experts verified questions for the languages of their expertise. The remaining $17$ participants were recruited from our institution and reviewed English questions for the regions they had lived in. Before beginning the task, all participants were provided with clear written instructions outlining the evaluation criteria, available in Appendix G.
Aya32B Qwen3-32B Gemma3-27B GPT4.1 LLaMA3.3 Gemini2.5Pro
@@
Table 9. Accuracy of six models in answering cultural knowledge questions on the full question bank (full dataset) and on questions derived from misrepresentations generated by that model (model subset). Best performance in each row is bold
Each question was independently assessed along four dimensions using a yes/no format: (1) validity: whether the question was valid; (2) uniqueness: whether it admitted a single unambiguous answer; (3) cultural grounding: verifying the presence of a cultural element; and (4) answer accuracy: confirming that the expected response was correct. In cases where a question did not satisfy one or more of these criteria, participants were instructed to refine it to meet the required standards. If the question received a ‘no’ on any of the dimensions and the verifier did not suggest a refinement, the question was excluded fromTALES-QA. This protocol ensured that the final set of questions included in TALES-QA were clear, accurate, and culturally appropriate.
The question generation and verification process resulted in TALES-QA, a question bank comprising $5$ different types of questions: one-word answer, one-phrase answer, fill-in-the-blank, multiple-choice, and true/false. The final bank contained a total of 1,683 questions, with 1,115 in Indic languages and 568 in English (see Table 8). Most of the questions are of one-word type, which are particularly effective at testing cultural knowledge. Unlike True/False and MCQs, one-word questions offer a more robust means of assessment because they reduce the possibility of random guessing.
5.2. Evaluation Methodology
We evaluated all $6$ models on TALES-QA to assess their cultural knowledge. To prevent models from producing long-form responses, we provided instructions tailored to each question type. To mitigate potential false negatives arising from Unicode or script matching issues in Indic languages, and following the recent success of auto raters for answer matching (Chandak et al., 2025), we used GPT 44o as a judge to compare each model’s responses against the correct answers (prompt provided in Appendix H). Additionally, to capture variability and uncertainty in model outputs, we sampled five responses from the model and computed average across all of them. We measured each model’s accuracy on the entire question bank. To further investigate whether misrepresentations were primarily due to a lack of relevant cultural knowledge, we also measured each model’s accuracy on the subset of questions derived specifically from misrepresentations present in the stories generated from that same model.
@@ 5.3. Findings
We also observed that for the closed-source models, GPT $4.1$ and Gemini 2.52.5Pro, accuracy decreased on the model-specific subsets compared to their performance on the overall English question bank. In contrast, for the open-source models, Gemma 33 2727B and LLaMa $3.3$ 7070B, accuracy increased on the model-specific subsets. This suggests that the questions derived from closed-source models may be more challenging than those derived from open-source models. Moreover, even for the best models, there exists headroom for improvement (from 86.386.3%) in answering cultural knowledge questions, in both English and Indic languages. We believe that TALES-QA will be a useful resource for testing cultural knowledge from India, in English and $13$ Indic languages, for future models.
@@ 6.1. Community-Centred Evaluations of Cultural Representation
Evaluation of an AI system, for any capability, requires a number of assumptions and design decisions (Jacobs and Wallach, 2021; Wallach et al., 2024). These assumptions and decisions are often implicit, almost exclusively made by developers of AI systems, and reflect the priorities and incentive structures embedded within AI development. Yet work has shown that values prioritized by model developers and the general public may not align (Jakesch et al., 2022) and what different communities consider desirable is subjective (Diaz and Madaio, 2023). This issue is likely to be exacerbated when AI developers do not represent the heterogeneity of the users and disproportionately reflect privileged sociocultural identities (Sambasivan et al., 2021). For evaluating cultural representation, such misalignment can be problematic because needs and desires of cultural representations are deeply intertwined with lived experiences of communities. Scholarship in HCI has long advocated for integrating values and perspectives of diverse stakeholders through frameworks like value sensitive and participatory design (Friedman and Hendry, 2019; Zytko et al., 2022) when evaluation decisions directly impact, and can potentially harm minoritized communities.
Keeping this in mind, we ground our evaluation of cultural misrepresentations by LLMs in the deep cultural understanding of community members. From the outset, we engage with participants with diverse cultural identities and lived experiences from across India to create a taxonomy of cultural misrepresentations in LLM-generated stories followed by a larger-scale human evaluation by expert annotators. This allowed us to combine scale and nuance in conducting fine-grained analyses of cultural misrepresentations leading to valuable insights that illuminate future research directions, including: (1) gaps in multilingual capabilities and cultural representation for peri-urban regions, (2) highlighting the cultural entities (practices, norms, and food) that are often misrepresented, and (3) demonstrating the gap between models’ cultural knowledge in answering standalone questions and ability to use this knowledge in generative settings. Finally, the taxonomy of cultural misrepresentations and questions of cultural knowledge for diverse Indian identities, will be useful for future research in this area.
@@ 6.2. Minoritized Communities within Minoritized Communities
Despite AI systems being deployed globally, they continue to better represent and serve some identities (Durmus et al., 2023; AlKhamissi et al., 2024). The studies of harms or cultural competence are often presented as a dichotomy between “western” and “non-western” contexts (Qadri et al., 2025). However, this formulation presumes a monolithic cultural identity of the minoritized “non-western” region. India is home to people from diverse linguistic, ethnic, religious, socioeconomic, and educational backgrounds. The axes of disparities among these communities are distinct from the rest of the world (Bhatt et al., 2022). Treating the diverse cultural identities within any (minoritized) region, as a homogeneous group is likely to prevent a finer-grained understanding of how AI systems may (mis)represent or harm minoritized communities within this region.
Our work echoes and adds to these findings: cultural misrepresentations by LLMs in the Indian context are not only prevalent, but they are exacerbated in lower-resourced Indic languages and for identities from less well-known regions (§4.3). This has particularly concerning implications for Indian users of digital and AI technology. As of 2021, India has the second-highest number of active internet users in the world (World Bank, 2025a; Wikipedia contributors, 2025). Over 60% of the population resides in rural areas (World Bank, 2025b) and 50% of Indian users prefer to access internet in Indic languages (AFPR, 2025; Statista, 2025). This underscores the urgent need for building technologies that engage with the lived experiences of diverse communities in their languages.
Building language technologies in the Indian context has received attention on various fronts, like curating pre-training corpora (Khan et al., 2024), developing benchmarks (Verma et al., 2025; Singh et al., 2024; Dev et al., 2023; Seth et al., 2024), building foundation models (Kallappa et al., 2025; Choudhury et al., 2025), evaluating fairness (Bhatt et al., 2022; Khandelwal et al., 2024) and cultural competence (Maji et al., 2025). We add to this by evaluating cultural representation in open-ended generation through engagement with people from a variety of regional, linguistic, and cultural backgrounds.
@@ 6.3. Designing Systems for Failure Modes
We find that LLMs fall in one of two failure modes: (1) generating generic text which does not include any specific cultural elements (no-representation), or (2) hallucinating information about the culture and generating text that is rife with inaccuracies around cultural elements (misrepresentations). Specifically, our findings (§4.3) indicate that while Llama3.3-70B generated fewer misrepresentations per story, it also had low CSI counts in the stories to begin with. This indicates that Llama3.3-70B might be generating generic stories that do not incorporate enough cultural nuance. On the other hand, Qwen3-32B and Aya-32B generate high number of CSIs in the stories but also have a larger number of misrepresentations per story. In other words, we could think of Llama3.3-70B is prone to the failure mode of no-representation, while Qwen3-32B and Aya-32B are prone to misrepresentation. Neither misrepresentation nor no-representation is an ideal outcome. In fact, the participants we engaged with indicated that the result of both these failure modes —be it inaccuracies in cultural elements or oversimplification of cultural elements—is dissatisfying.
We argue that AI systems must be thoughtfully designed to handle situations where the system encounters requests requiring cultural representations that may be under-specified, subjective, cannot be fulfilled based on the parametric or external knowledge that the system can access. AI systems may be designed to handle such situations in a variety of ways to prevent giving wrong or misleading answers. For example, the LLM may be designed to ask follow-up questions (Li et al., 2024a, 2025), abstain from answering (Wen et al., 2025; Brahman et al., 2024), or provide multiple possible outputs (Sorensen et al., 2024). Moreover, such a design must also account for the sociocultural context of the deployment, since users across cultures may have varying expectations (Fortunati et al., 2022; Ge et al., 2024) and in turn varying tolerance to different model behaviours (Lucy et al., 2023). As AI systems become integrated into people’s daily lives, future research must investigate the impacts of their design and behavior on the experiences and the harms faced by diverse users.
@@ 7. Limitations and Future Work
Although we view this work as an important step towards understanding and evaluating misrepresentation in generated stories, we acknowledge a few important limitations of our work that should be investigated in future research. First, all focus groups and surveys were conducted on English-language stories generated by GPT-$4$, primarily to facilitate consistent interpretation and discussion within the research team. Further research is required to assess whether any new categories would emerge from other languages or from other models. Second, our evaluation study focused primarily on cultural contexts within India, and thus the findings were shaped by the perspectives most familiar to our expert annotators. While this still provides valuable insights, future work could utilize our methodological framework to expand data collection beyond India to capture a broader range of cultural contexts and to examine how similar or dissimilar misrepresentations manifest globally. Finally, in line with prior work (Bhatt and Diaz, 2024; Bhagat et al., 2025), we used open-ended prompts that describe the story topic and cultural identity and evaluate cultural representations in these narratives. Specifying only the identity, without additional details about the cultural nuances was aimed to elicit the model’s implicit cultural representations rather than relying on information supplied in the prompt. This prompting design may have impacted the LLMs’ ability to accurately and appropriately represent cultural identities, and therefore, the proportion of misrepresentations observed may change under more specific or richer prompts. While our taxonomy is grounded in the prompts we used, we believe it offers value for evaluating cultural misrepresentations across a range of prompting strategies. LLMs are fast-evolving, and our work provides a useful starting point rather than a definitive endpoint; future work could explore the impact of more detailed prompts, multi-turn interactions, or external knowledge sources on the models’ cultural representations.
Our work also points to concrete research directions to improve LLMs’ cultural competence for non-Western contexts in generative settings. Specifically, our taxonomy provides a structured way to identify and analyze misrepresentations and future work should build on this to develop concrete strategies for mitigating the cultural misrepresentations we observed. Finally, future work must investigate why LLMs generate misrepresentations in open-ended generations despite possessing the requisite cultural knowledge.
@@ 8. Conclusion
Acknowledgements.
We are grateful to the participants of our focus groups and survey, and the expert annotators of our study for their time, expertise, and thoughtful engagement with our study. We thank folks at AI4Bharat, specifically, Mitesh Khapra and Sounak Dutta, for facilitating the annotation process. This work was supported in part by the AI2050 program at Schmidt Sciences (Grant G-24-66186) and a grant from Google Research (thanks to Dinesh Tewari). We are grateful to Kinshuk Vasisht for his feedback and for being the external coder for the development of our taxonomy. We thank Jimin Mun, Akhila Yerukola, Joel Mire, Mansi Gupta, Lora Aroyo, Partha Talukdar and Deepak Varuvel Dennison for their feedback on the manuscript. Lastly, we are grateful to Vani A. for her help with project coordination and administration.
@@ Appendix A Demographic details for focus groups and surveys
For detailed demographic information, please refer to Table 10
@@ Appendix B Questions Used by Facilitator in Focus Groups and Surveys
      What do they look for in a story? Is relatability important?
@@
  • Questions after participants have completed reading stories:
@@
      Are there any similarities between the stories?
Focus Group 1: Jodhpur, Ahmedabad, Jaipur
(Gujarat and Rajasthan)
Focus Group 2: Kannur, Mayiladuthurai, Chennai
(Kerala and Tamil Nadu)
Focus Group 3: Bhubaneswar, Kolkata, Bargarh
(Odisha and West Bengal)